OpenStudy (anonymous):

find the axis of symmetry and vertex of x squared + 5x- 12

OpenStudy (anonymous):

for any parabola in the form: \[f(x) = ax ^{2} = bx + c\] you can find the axis of symmetry using using the equation below: \[x = -b \div 2a\] that then gives you your x coordinate which is the axis of symmetry (note that this is also the x coordinate for the vertex, by substituting the x value back in to the original equation you will get the y coordinate of the vertex). therefore the above equation will equal: \[x=-5\div2(12)\] \[= -5/24\]

OpenStudy (anonymous):

f(x)=ax2+bx+c*** fixes the + sign to a = sign which is an error in the first equation